A borrower arranged a 250,000,000 credit line with a bank.  It drew down:

  • 20% for the first quarter of the year
  • An additional 20% for the second quarter
  • The remaining for the second half.

            Assume the interest rate was 3% for the first quarter, and 3.2% for the rest of the year.

            The unused credit line fee was 0.2%.

            Ignore day counts for this question; that is, just use ¼ for each three month period.

  1. What is the interest amount and fee amount for each quarter?
  2. What is the effective “all in” cost of the borrowing?  Do this by tallying up the average amount actually borrowed over the course of the year.  Then calculate the total amount paid to the bank, and divide.
Answer #1

a) calculation of interest amount:

I quarter: 3*(1/4)% of (20% of 250000000) = 375000

II quarter: 3.2(1/4)% of (40% of 250000000) = 800000

III quarter: 3.2(1/4)% of (100% of 250000000) = 2000000

IV quarter: 3.2(1/4)% of (100% of 250000000) = 2000000

Total interest amount = 5175000

calculation of fee amount:

I quarter: .2*(1/4)% of (80% of 250000000) = 100000

II quarter: .2(1/4)% of (60% of 250000000) = 75000

III quarter: nil

IV quarter: nil

Total fee amount = 175000

b) calculation of average amount borrowed =

250000000*.2*(12/12) + 250000000*.2(9/12) + 250000000*.6(6/12)

=162500000

Totat interest and fee amount = 5175000 + 175000 = 5350000

Effective cost = 5350000/162500000

= 3.29%